Let X be the number of life years before certain kind of CFL bulb needs replacement . Let X have the probability mass function is defined as

f(x)=alphax**3,x=0,1,2,3,4

Where alpha is constant

(a) Find the values of constant alpha

(b) find the cumulative distribution function F to X



1
Expert's answer
2022-03-07T11:13:07-0500

(a)


"\alpha(0)^3+\alpha(1)^3+\alpha(2)^3+\alpha(3)^3+\alpha(4)^3=1"

"\alpha=0.01"


(b)


"f(x)=\begin{cases}
 0 &x=0\\
 0.01&x=1 \\
 0.08 &x=2 \\
 0.27 & x=3 \\
 0.64 & x=4\\
\end{cases}"

Then

"F(x)=\begin{cases}
 0 &x<1\\
 0.01&1\le x<2 \\
 0.09 &2\le x<3 \\
 0.36 & 3\le x<4 \\
 1 & x\ge 4 \\
\end{cases}"